OpenStudy (anonymous):

Help please? :o 1. Solve the system by substitution- 2x+y=-11, 3x-4y=11. I'm so terrible at substitution and elimination >.< 2. Solve the system using elimination- 2x+6y=-12, 5x-5y=10. 3.What is the solution of the following system?-3x-2y=-12, 9x+6y=-9. Help with any of these would be greatly appreciated thanks.

OpenStudy (anonymous):

Even any tips on how to do these would be nice :/

OpenStudy (anonymous):

ok so looking at the first one. basically you have to multiply one or both of the equations by a number so that the coefficient of one of the variables is the same so that when you subtract/add them then it will cancel. ^that sounds confusing, here's an ex for elimination: x+y=5, 2x+y = 30 so if we multiplied the first equation by 2, we'd get: 2x+2y=10 right? so we can subtract now... 2x+2y=10 - 2x+y=30 ------------ 0 + y = -20 ------------- and now we can plug in -20 for y back in the equation to find x. hope that made sense

OpenStudy (anonymous):

Yeah that made more sense than my textbook does haha. How about a substitution example? Or how you can tell if a system is dependant or independant? or has infinately many solutions?

OpenStudy (anonymous):

haha yeah textbooks can be annoying sometimes. so substitution: x+y=2 x+4y = 7 so in substitution we basically just have to get any variable on its own so that we can substitute it into another equation. pretty much cause solving for 2 variables is hard. so i can get either x or y by itself from the 1st equation (or the second if you want). x+y = 2 so x+y-y = 2-y so, x+0=2-y or x=2-y now we know what x equals, so lets put it in the other equation. (x) +4y=7 OR (2-y) +4y = 7 [see what i did there? ] so 2-y+4y=7, 2+3y=7, 3y=5, y=5/3 and you can put this in the x equation (x=2-y) to get x aswell :)

OpenStudy (anonymous):

Very much so >.< Man that all still seems pretty confusing but I really appreciate the help! I always just tend to get one step wrong along the way and mess the whole thing up lol, It's in my nature :P

OpenStudy (anonymous):

x-y=11, -x+y=-11 Would one like this have infinitely many solutions?

OpenStudy (anonymous):

same here. i used to hate math so much, i still sort of suck at it. even if i get the stuff i make little mistakes. <.< anyway, if you just practice it a bit i'm sure you'll get it. yeah that would have infinitely many solutions because they're the same equation just written differently. so on a graph, there would be just the one line. this might help: http://www.mathocean.com/2009/10/system-of-linear-equations-consistency.html

OpenStudy (anonymous):

Well I'll keep trying x.x I just need to do decently on my daily tests which so far I haven't really haha. And yeah that's what i though lol. Funny I swear I've been to that website before x)

OpenStudy (anonymous):

daily tests? ai yi yi. haha well that's good, good luck!

OpenStudy (anonymous):

Yeah every day, about 10 question tests x.x it's ridiculous.. But thanks !

OpenStudy (anonymous):

woah that is a little ridiculous. well you get practice from it i suppose. yeah no problem :D